# Dual thermal pseudo-critical features in a spin-1/2 Ising chain with twin-diamond geometry

### v1

@@ Abstract
One-dimensional lattice models have recently attracted renewed attention because certain systems display thermodynamic anomalies that closely resemble finite-temperature phase transitions [1, 2, 3]. These anomalies, known as pseudo-transitions, differ from genuine criticality even though they reproduce several of its characteristic signatures. First derivatives of the free energy, such as entropy and magnetization, exhibit steep but continuous variations near a pseudo-critical temperature, while second derivatives, including the specific heat and magnetic susceptibility, develop sharp yet finite peaks [4]. Numerical and analytical studies of decorated Ising and Ising-Heisenberg chains have further shown that these pseudo-transitions can mimic essential features of critical phenomena: the correlation length increases rapidly near the pseudo-critical temperature [5], and several response functions follow power-law forms governed by a universal set of pseudo-critical exponents identified in many one-dimensional decorated models [6].
@@
Peculiar decorated one-dimensional spin models provide a natural framework for these anomalies because their internal degrees of freedom generate competing low-energy sectors whose exchange of dominance at a characteristic temperature produces fully analytic yet remarkably sharp thermodynamic crossovers. This mechanism has been identified in several exactly solvable systems, including the Ising-XYZ diamond chain, where quasi-phases and pseudo-transitions were first clearly characterized [4, 5], the coupled spin-electron double-tetrahedral chain, which exhibits frustration-enhanced entropy plateaus [1], and various Ising-Heisenberg ladders and tubes that display pseudo-critical behavior governed by universal exponents [2, 6, 7, 8, 9, 10]. A unified interpretation based on the decoration-iteration mapping clarified that the anomalous crossover originates from temperature-dependent effective couplings and effective fields generated after integrating out the decorated sublattices [6]. Furthermore, subsequent analyses established that the residual entropy at zero-temperature phase boundaries serves as a reliable predictor for the appearance of pseudo-transitions in decorated chains [11].
@@
Pseudo-transition behavior is not restricted to decorated chains. Several strictly one-dimensional models without auxiliary spins or internal bond structures also display sharp thermal crossovers when their parameters are tuned to generate competing low-energy configurations. A notable example is the extended Hubbard model at half filling in the atomic limit, where pronounced thermodynamic anomalies mimic those associated with quasi-phases [12]. A similar crossover between antiferromagnetic and charge-ordered configurations was reported in a minimal spin-pseudospin description of cuprate chains, accompanied by a strong but finite specific heat peak [13]. Related effects occur in frustrated Potts-type models such as the q-state Potts-Zimm-Bragg formulation [14], in dilute Ising chains with impurity induced frustration [15], and in water-like lattice models incorporating van der Waals interactions and directional bonding [16]. These examples demonstrate that the essential ingredients behind pseudo-transitions, namely competing sectors with small energy gaps and enhanced degeneracies, can arise naturally even in purely one-dimensional systems without explicit decoration.
@@
Diamond chain geometries play a central role in several magnetic compounds, which explains why diamond chain models have been extensively studied in the theoretical literature [3, 17, 18, 19]. A prominent example is the natural mineral azurite $\mathrm{Cu_{3}(CO_{3})_{2}(OH)_{2}}$[20, 21], whose magnetic properties motivated detailed analyses of spin 1/2 Ising and Ising-Heisenberg diamond chains. Other compounds, such as the mixed valent iron diamond chain with single ion anisotropy [22], further support the idea that the diamond motif is an efficient source of frustration and anomalous thermodynamic behavior in one dimensional systems, among other similar structures[7, 23, 24]. For these reasons, diamond chains remain key reference models for exploring the origin of entropy plateaus, degeneracy driven anomalies, and pseudo-critical behavior in one dimension.
@@
Beyond simple diamond chains, an increasing number of materials display coupled diamond like building blocks, often arranged in layers or extended structural networks. A representative case is the oxohalide $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$, where $\mathrm{Cu}^{2+}$ ions form diamond shaped units that are connected along a crystallographic direction [25]. Similar motifs appear in more complex families such as the $\mathrm{CuO\text{\textendash}CuCl_{2}\text{\textendash}SeO_{2}}$ series [26], as well as in the oxohalides $\mathrm{Cu}_{3}(\mathrm{SeO}_{3})_{2}\mathrm{Cl}_{2}$ and $\mathrm{Cu}_{3}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$ [27], and the insulating spin-gap compound $\mathrm{CdCu_{2}(SeO_{3})_{2}Cl_{2}}$ [28]. Such structural features strongly motivate the study of theoretical models that incorporate more than one diamond unit per cell. The coupled twin diamond chain examined in the present work represents the minimal extension capable of capturing this additional connectivity and allows us to investigate how the interaction between two diamond units within each cell can produce multiple frustrated sectors and distinct pseudo-critical features in a single exactly solvable framework.
@@
This work is organized as follows. Section II introduces the Hamiltonian, outlines the transfer-matrix solution, and presents the zero-temperature phase diagram together with the corresponding pseudo-transition features. Section III examines the thermodynamic behavior in the anomalous region where these pseudo-transitions emerge. Finally, Section IV summarizes the main conclusions.
@@ II Coupled twin-diamond chain model
In this section we consider a coupled twin-diamond chain (CTDC) inspired by the compound ${\rm Cu}_{2}({\rm TeO}_{3})_{2}{\rm Br}_{2}$ whose magnetic structure consists of spin-1/2 ${\rm Cu}^{2+}$ ions arranged in diamond-like units[25]. Here we assume the simplest case that the magnetic interactions are purely of the Ising type. A schematic representation of the lattice geometry is shown in Fig. 1. Blue spheres denote the nodal spins, whereas black spheres indicate the two spins forming each internal dimer.
@@
Figure 1: (Top) Schematic representation of the spin-1/2 CTDC. Black spheres denote dimer spins ($S_{a,k}$ and $S_{b,k}$), while blue spheres represent nodal Ising spins. Blue (red) lines indicate the couplings between dimer and nodal spins, associated with the exchange interactions $J_{1}$ ($J_{0}$), respectively, whereas the black line corresponds to the intra-dimer exchange interaction ($J_{2}$). (Middle) Illustration of the decoration transformation. (Bottom) Resulting effective spin chain with nearest- and next-nearest-neighbor interactions.
@@ II.1 Hamiltonian and transfer matrix
Here $J_{2}$ is the intra-dimer coupling, $J_{0}$ couples the nodal spin $s_{k}$ to the adjacent dimer, and $J_{1}$ couples the dimer to the neighboring nodal spins $s_{k-1}$ and $s_{k+1}$ along the chain. Due to $g_{0}\neq g_{1}$ in general, the two sublattices respond differently to the same external magnetic field $B$. The Hamiltonian (1) is invariant under the exchange $S_{a,k}\leftrightarrow S_{b,k}$ and under cyclic permutations of the three nodal spins $s_{k-1},s_{k}$ and $s_{k+1}$.
To investigate the thermodynamic properties we employ the transfer-matrix method, which provides an exact treatment for one-dimensional Ising-type models. In the present case a decoration-iteration transformation[29, 30, 31] maps the CTDC onto an effective Ising chain with next-nearest-neighbor interactions. Consequently, the transfer matrix $V$ can be written in a $4\times 4$ form, analogous to that of the extended Ising chain with next-nearest couplings [32],
$V=\left(\begin{array}[]{cccc}\mathfrak{r}&\mathfrak{m}&0&0\\ 0&0&\mathfrak{p}&\mathfrak{n}\\ \mathfrak{m}&\mathfrak{q}&0&0\\ 0&0&\mathfrak{n}&\mathfrak{t}\end{array}\right),$ (2)
where the basis is labeled by the pair ($s_{k},s_{k+1}$). The matrix elements $\mathfrak{r}$, $\mathfrak{m}$, $\mathfrak{p}$, $\mathfrak{n}$, $\mathfrak{q}$, $\mathfrak{t}$ incorporate the local Boltzmann weights for the four possible configurations of neighboring nodal spins and the internal dimer
@@
$a_{2}=$ $-3\frac{\mathfrak{c}_{3}^{2}}{8}+\mathfrak{c}_{2}.$
@@
where the auxiliary quantities are
$y=$ $\frac{a_{2}}{6}+w-\frac{p}{3w}$ (19)
$w=$ $\sqrt[3]{-\frac{q}{2}+\sqrt{\frac{q^{2}}{4}+\frac{p^{3}}{27}}},$ (20)
$p=$ $-\frac{a_{2}^{2}}{12}-a_{0},$ (21)
$q=$ $-\frac{a_{2}^{3}}{108}+\frac{a_{2}a_{0}}{3}-\frac{a_{1}^{2}}{8}.$ (22)
@@ III Zero-temperature phases and phase diagram
We adopt the single-cell basis $|S_{a,k},S_{b,k};s_{k}\rangle$, with each spin variable $S_{a,k},S_{b,k},s_{k}=\pm\tfrac{1}{2}$. For a chain of $N$ unit cells the Hamiltonian decomposes into local blocks, and the ground-state configurations of these blocks determine the zero-temperature phase diagram. The low-temperature entropy map in Fig. 2 provides a direct visualization of the resulting phases: nonfrustrated regions appear in cyan, with vanishing entropy, while frustrated sectors form green and orange plateaus with $S/k_{\mathrm{B}}\simeq\tfrac{1}{2}\ln(2)$ and $S/k_{\mathrm{B}}\simeq\ln(2)$, respectively.
@@ III.1.3 Fully frustrated phase
The second frustrated ($\mathrm{FR}_{2}$) occurs when all nodal spins are polarized upward, $s_{k}=\tfrac{1}{2}$, while the dimer spins remain antiparallel. A convenient representative ground state is
@@ III.1.5 Antiferromagnetic phase
Taken together, the five phases SP, FI, AF, $\mathrm{FR}_{1}$, and $\mathrm{FR}_{2}$ account for all ground-state regions observed in Fig. 2. The frustrated sectors produce the characteristic entropy plateaus, while the nondegenerate ordered regiosn is identified as low-entropy (cyan regions). Their boundaries constitute the zero-temperature skeleton underlying the sharp but continuous low-temperature crossovers discussed in the next subsection.
@@ III.2 Phase boundary
III.3 Pseudo-critical temperature
A pseudo-critical temperature $T_{p}$ along any fixed parameter cut, for instance at fixed $J_{0}$, $J_{1}$, $J_{2}$, while varying $B$, follows from the crossing of the low temperature free energy branches of the two competing phases. Approximating each free energy by $f_{A}(T)\simeq\varepsilon_{A}-Ts_{A}$, the condition $f_{A}(T_{\mathrm{p}})=f_{B}(T_{\mathrm{p}})$ yields the standard estimate of pseudo-critical temperature

### v2

@@ Abstract
One-dimensional lattice models have recently attracted renewed attention because certain systems display thermodynamic anomalies that closely resemble finite-temperature phase transitions (galisova, ; strecka16, ; sm-souza16, ). These anomalies, known as pseudo-transitions, differ from genuine criticality even though they reproduce several of its characteristic signatures. First derivatives of the free energy, such as entropy and magnetization, exhibit steep but continuous variations near a pseudo-critical temperature, while second derivatives, including the specific heat and magnetic susceptibility, develop sharp yet finite peaks (sergio18, ). Numerical and analytical studies of decorated Ising and Ising-Heisenberg chains have further shown that these pseudo-transitions can mimic essential features of critical phenomena: the correlation length increases rapidly near the pseudo-critical temperature (Icarvalho, ), and several response functions follow power-law forms governed by a universal set of pseudo-critical exponents identified in many one-dimensional decorated models (Rojas2019, ).
@@
Peculiar decorated one-dimensional spin models provide a natural framework for these anomalies because their internal degrees of freedom generate competing low-energy sectors whose exchange of dominance at a characteristic temperature produces fully analytic yet remarkably sharp thermodynamic crossovers. This mechanism has been identified in several exactly solvable systems, including the Ising-XYZ diamond chain, where quasi-phases and pseudo-transitions were first clearly characterized (sergio18, ; Icarvalho, ), the coupled spin-electron double-tetrahedral chain, which exhibits frustration-enhanced entropy plateaus (galisova, ), and various Ising-Heisenberg ladders and tubes that display pseudo-critical behavior governed by universal exponents (strecka16, ; Rojas2019, ; Derzhko, ; dmd-potts, ; Krokhmalskii, ; Chapman, ). A unified interpretation based on the decoration-iteration mapping clarified that the anomalous crossover originates from temperature-dependent effective couplings and effective fields generated after integrating out the decorated sublattices (Rojas2019, ). Furthermore, subsequent analyses established that the residual entropy at zero-temperature phase boundaries serves as a reliable predictor for the appearance of pseudo-transitions in decorated chains (bjp20, ).
@@
Pseudo-transition behavior is not restricted to decorated chains. Several strictly one-dimensional models without auxiliary spins or internal bond structures also display sharp thermal crossovers when their parameters are tuned to generate competing low-energy configurations. A notable example is the extended Hubbard model at half filling in the atomic limit, where pronounced thermodynamic anomalies mimic those associated with quasi-phases (verissimo, ). A similar crossover between antiferromagnetic and charge-ordered configurations was reported in a minimal spin-pseudospin description of cuprate chains, accompanied by a strong but finite specific heat peak (katarina, ). Related effects occur in frustrated Potts-type models such as the q-state Potts-Zimm-Bragg formulation (Panov2021, ), in dilute Ising chains with impurity induced frustration (Yasinskaya, ), and in water-like lattice models incorporating van der Waals interactions and directional bonding (Braz25, ). These examples demonstrate that the essential ingredients behind pseudo-transitions, namely competing sectors with small energy gaps and enhanced degeneracies, can arise naturally even in purely one-dimensional systems without explicit decoration.
@@
Diamond chain geometries play a central role in several magnetic compounds, which explains why diamond chain models have been extensively studied in the theoretical literature (sm-souza16, ; strecka-dmd, 2020, 2004, ). A prominent example is the natural mineral azurite $\mathrm{Cu_{3}(CO_{3})_{2}(OH)_{2}}$(kikuchi03, ; kikuchi04, ), whose magnetic properties motivated detailed analyses of spin 1/2 Ising and Ising-Heisenberg diamond chains. Other compounds, such as the mixed valent iron diamond chain with single ion anisotropy (Sorolla, ), further support the idea that the diamond motif is an efficient source of frustration and anomalous thermodynamic behavior in one dimensional systems, among other similar structures(Derzhko, ; zad, ; Chakhmakhchyan, ). For these reasons, diamond chains remain key reference models for exploring the origin of entropy plateaus, degeneracy driven anomalies, and pseudo-critical behavior in one dimension.
@@
Beyond simple diamond chains, an increasing number of materials display coupled diamond like building blocks, often arranged in layers or extended structural networks. A representative case is the oxohalide $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$, where $\mathrm{Cu}^{2+}$ ions form diamond shaped units that are connected along a crystallographic direction (Uematsu, 2007). Similar motifs appear in more complex families such as the $\mathrm{CuO\text{\textendash}CuCl_{2}\text{\textendash}SeO_{2}}$ series (Kakarla, ), as well as in the oxohalides $\mathrm{Cu}_{3}(\mathrm{SeO}_{3})_{2}\mathrm{Cl}_{2}$ and $\mathrm{Cu}_{3}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$ (Rocquefelte, ), and the insulating spin-gap compound $\mathrm{CdCu_{2}(SeO_{3})_{2}Cl_{2}}$ (Murtazoev, ). Such structural features strongly motivate the study of theoretical models that incorporate more than one diamond unit per cell. The coupled twin diamond chain examined in the present work represents the minimal extension capable of capturing this additional connectivity and allows us to investigate how the interaction between two diamond units within each cell can produce multiple frustrated sectors and distinct pseudo-critical features in a single exactly solvable framework.
In the present work, the coupled twin-diamond chain is therefore introduced as a minimal theoretical model inspired by such diamond-based magnetic structures(Uematsu, 2007), rather than as a quantitative description of any specific compound. The central novelty of the coupled twin-diamond chain lies in the fact that its minimal unit cell supports more than one frustrated low-energy manifold with different extensive degeneracies. As a result, the system exhibits two well-separated pseudo-critical temperature scales associated with successive entropy-driven crossovers between competing configurations. This provides a clean exactly solvable example in which dual pseudo-transition features arise within a single one-dimensional decorated Ising geometry, rather than from tuning parameters between distinct models. We further show that these features remain robust even when the two spin sublattices respond differently to an external magnetic field.
@@
This work is organized as follows. Sect. II introduces the Hamiltonian, outlines the transfer-matrix solution, and presents the zero-temperature phase diagram together with the corresponding pseudo-transition features. Sect. III examines the thermodynamic behavior in the anomalous region where these pseudo-transitions emerge. Finally, Sec. IV summarizes the main conclusions.
@@ II Coupled twin-diamond chain model
In this section we consider a coupled twin-diamond chain (CTDC) inspired by the compound ${\rm Cu}_{2}({\rm TeO}_{3})_{2}{\rm Br}_{2}$ whose magnetic structure consists of spin-1/2 ${\rm Cu}^{2+}$ ions arranged in diamond-like units(Uematsu, 2007). Here we assume the simplest case that the magnetic interactions are purely of the Ising type. A schematic representation of the lattice geometry is shown in Fig. 1. Blue spheres denote the nodal spins, whereas black spheres indicate the two spins forming each internal dimer.
@@
Figure 1: (Top) Schematic representation of the spin-1/2 CTDC. Black spheres denote dimer spins ($S_{a,k}$ and $S_{b,k}$), while blue spheres represent nodal Ising spins. Blue (red) lines indicate the couplings between dimer and nodal spins, associated with the exchange interactions $J_{1}$ ($J_{0}$), respectively, whereas the black line corresponds to the intra-dimer exchange interaction ($J_{2}$). (Middle) Illustration of the star-triangle transformation. (Bottom) Resulting effective spin chain with nearest- and next-nearest-neighbor interactions.
@@ II.1 Hamiltonian and transfer matrix
Here $J_{2}$ is the intra-dimer coupling, $J_{0}$ couples the nodal spin $s_{k}$ to the adjacent dimer, and $J_{1}$ couples the dimer to the neighboring nodal spins $s_{k-1}$ and $s_{k+1}$ along the chain. Due to $g_{0}\neq g_{1}$ in general, the two sublattices respond differently to the same external magnetic field $B$. The Hamiltonian (1) is invariant under the exchange $S_{a,k}\leftrightarrow S_{b,k}$ and is translationally invariant along the chain.
At this stage, the internal dimer spins $S_{a,k}$ and $S_{b,k}$ can be eliminated by means of a local star-triangle transformation (dec-trns, ; ddec-trns, ; jozef-dctrns, ), as schematically illustrated in Fig. 1 (middle). For fixed values of the neighboring nodal spins, the Boltzmann weight of each decorated cell is obtained by explicitly summing over the dimer degrees of freedom. The repeated application of this mapping along the chain yields an effective one-dimensional Ising model for the nodal spins, with temperature-dependent nearest- and next-nearest-neighbor interactions and an effective field.
[…]
$\mathbf{V}=\left(\begin{array}[]{cccc}\mathfrak{r}&\mathfrak{m}&0&0\\ 0&0&\mathfrak{p}&\mathfrak{n}\\ \mathfrak{m}&\mathfrak{q}&0&0\\ 0&0&\mathfrak{n}&\mathfrak{t}\end{array}\right),$ (2)
where the basis states are labeled by the nodal-spin pairs ($s_{k},s_{k+1}$). The matrix elements $\mathfrak{r}$, $\mathfrak{m}$, $\mathfrak{p}$, $\mathfrak{n}$, $\mathfrak{q}$, $\mathfrak{t}$ incorporate the local Boltzmann weights for the four possible configurations of neighboring nodal spins and the internal dimer
@@
$a_{2}=$ $-3\frac{\mathfrak{c}_{3}^{2}}{8}+\mathfrak{c}_{2}<0.$
@@
where the $y$ is choosing for conveniency as
$y=$ $2\sqrt{Q}\cos\left(\tfrac{R}{3\sqrt{Q^{3}}}\right)+\frac{a_{2}}{6},$ (19)
with
$Q=$ $\frac{a_{2}^{2}}{36}+\frac{a_{0}}{3},$ (20)
$R=$ $\frac{a_{2}^{3}}{216}-\frac{a_{2}a_{0}}{6}+\frac{a_{1}^{2}}{16}.$ (21)
It is worth noting that the largest eigenvalue $\lambda_{1}$ is always real and strictly positive. On the branch associated with the dominant Perron(cuesta, ) eigenvalue, the second eigenvalue $\lambda_{2}$ is also real and satisfies $\lambda_{1}>\lambda_{2}$, as demonstrated in Appendix A. The remaining eigenvalues $\lambda_{3}$ and $\lambda_{4}$ may be either real or form a complex-conjugate pair, depending on the Hamiltonian parameters. Moreover, no general ordering can be inferred for the magnitudes $|\lambda_{2}|$, $|\lambda_{3}$|, and $|\lambda_{4}|$, which may vary across parameter regimes (see Appendix A).
@@ III Zero-temperature phases and phase diagram
We adopt the single-cell basis $|S_{a,k},S_{b,k};s_{k}\rangle$, with each spin variable $S_{a,k},S_{b,k},s_{k}=\pm\tfrac{1}{2}$. For a chain of $N$ unit cells the Hamiltonian decomposes into local blocks, and the ground-state configurations of these blocks determine the zero-temperature phase diagram. Frustrated ground-state manifolds may exhibit extensive degeneracy, leading to a finite residual entropy $\mathcal{S}_{0}\equiv\lim_{T\to 0}\mathcal{S}(T)$. At low but finite temperatures, this residual entropy manifests itself as nearly constant entropy plateaus. The low-temperature entropy map shown in Fig. 2 provides a direct visualization of this behavior: nonfrustrated regions appear in cyan, with vanishing entropy, while frustrated sectors form green and orange plateaus with $\mathcal{S}/k_{\mathrm{B}}\simeq\tfrac{1}{2}\ln 2$ and $\mathcal{S}/k_{\mathrm{B}}\simeq\ln 2$, respectively.
@@ III.1.3 Fully frustrated phase
The second frustrated ($\mathrm{FR}_{2}$) occurs when all nodal spins are polarized upward, $s_{k}=\tfrac{1}{2}$, while the dimer spins remain antiparallel. A convenient representative ground-state configuration is
@@ III.1.5 Antiferromagnetic phase
Taken together, the five phases SP, FI, AF, $\mathrm{FR}_{1}$, and $\mathrm{FR}_{2}$ account for all ground-state regions observed in Fig. 2. The frustrated sectors produce the characteristic entropy plateaus, while the nondegenerate ordered regions is identified as low-entropy (cyan regions). Their boundaries constitute the zero-temperature skeleton underlying the sharp but continuous low-temperature crossovers discussed in the next subsection.
@@ III.2 Phase boundary
III.3 Nearly block-diagonal form of the transfer matrix
Since the transfer matrix $\mathbf{V}$ given in Eq. (2) is irreducible, the Non-existence of phase transition theorem(cuesta, ) guarantees the absence of true phase transitions. Nevertheless, $\mathbf{V}$ may become nearly block diagonal in certain parameter regimes. In this case, the approach proposed in Ref. (gen-conds, ) provides a useful framework to analyze pseudo-critical behavior from a general perspective, without this tool it would become a cumbersome task find the pseudo-critical temperature. Following this approach, we reorganize the transfer matrix into a nearly block-diagonal form in order to identify the competing thermodynamic sectors responsible for the observed pseudo-critical temperatures.
III.3.1 Pair-sector decomposition
[…]
This basis can be conveniently splits into two distinct sectors:
(i) Parallel sector {$\{|++\rangle,\,|--\rangle\}$, corresponding to locally uniform spin configurations.
(ii) Antiparallel sector {$\{|+-\rangle,\,|-+\rangle\}$, corresponding to locally alternating nodal spin configurations.
In terms of the ground-state patterns discussed above, the $\mathrm{FR}_{2}$ and FI manifolds are predominantly supported on parallel sector, whereas the $\mathrm{FR}_{1}$ manifold is supported on antiparallel sector. Consequently, the $\mathrm{FI}-\mathrm{FR}_{1}$ and $\mathrm{FR}_{1}-\mathrm{FR}_{2}$ pseudo-transitions can be viewed as successive competitions between these two pair sectors.
III.3.2 Nearly block-diagonal structure of the transfer matrix
[…]
$\mathbf{V}=\begin{pmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{D}&\mathbf{B}\end{pmatrix}$, (30)
where the diagonal blocks are given by
$\mathbf{A}=\begin{pmatrix}\mathfrak{r}&0\\ 0&\mathfrak{t}\end{pmatrix}$, $\qquad\mathbf{B}=\begin{pmatrix}0&\mathfrak{p}\\ \mathfrak{q}&0\end{pmatrix}$, (31)
and the off diagonal blocks read
$\mathbf{C}=\begin{pmatrix}\mathfrak{m}&0\\ 0&\mathfrak{n}\end{pmatrix},\qquad\mathbf{D}=\begin{pmatrix}0&\mathfrak{n}\\ \mathfrak{m}&0\end{pmatrix}$. (32)
The diagonal blocks $\mathbf{A}$ and $\mathbf{B}$ propagate configurations within the parallel and antiparallel sectors, respectively, while the off-block matrices $\mathbf{C}$ and $\mathbf{D}$ encode processes that convert one sector into the other, which may be interpreted as domain-wall-like events.
Although $\mathbf{V}$ is irreducible, the inter-sector couplings $\mathbf{C}$ and $\mathbf{D}$ can become parametrically small in the low-temperature regime. In this sense, the transfer matrix is nearly block diagonal: the dominant statistical weight is carried by one sector at a time, and pseudo-critical behavior emerges when the leading eigenvalues associated with different sectors become nearly degenerate. This nearly reducible structure provides the spectral mechanism underlying the sharp yet analytic crossovers discussed in the main text.
The eigenvalues of the block $\mathbf{A}$ are trivially $\lambda=\{\mathfrak{r},\ \mathfrak{t}\},$ while the eigenvalues of the block $\mathbf{B}$ are $\lambda=\pm\sqrt{\mathfrak{p}\mathfrak{q}}$.
III.3.3 Pseudo-critical temperature
According to the criterion discussed in Ref. (gen-conds, ), pseudo-critical behavior arises when the transfer matrix acquires a nearly block-diagonal structure, such that the leading eigenvalues associated with competing blocks become comparable. In the present case, this leads to two possible pseudo-critical conditions, depending on which eigenvalue of the block $\mathbf{A}$ is dominant:
$\mathfrak{t}=\sqrt{\mathfrak{p}\mathfrak{q}}$ $\quad\Rightarrow\;T_{p_{1}}:\quad\text{associted to}\quad\mathrm{FR}_{1}-\mathrm{FI},$ (33)
$\mathfrak{r}=\sqrt{\mathfrak{p}\mathfrak{q}}$ $\quad\Rightarrow\;T_{p_{2}}:\quad\text{associted to}\quad\mathrm{FR}_{2}-\mathrm{FI}.$ (34)
These relations provide a transparent interpretation of the two pseudo-critical temperature scales observed in the thermodynamic analysis. It is worth emphasizing that they follow directly from the criterion of Ref. (gen-conds, ). Although simple in form, the resulting conditions involve transcendental equations.
Alternatively, the pseudo-critical temperature $T_{p}$ can also be determined along any fixed parameter cut in the low-temperature regime. For instance at fixed $J_{0}$, $J_{1}$, $J_{2}$, while varying $B$, follows from the crossing of the low temperature free energy branches of the two competing phases. Approximating each free energy by $f_{A}(T)\simeq\varepsilon_{A}-Ts_{A}$, the condition $f_{A}(T_{\mathrm{p}})=f_{B}(T_{\mathrm{p}})$ yields the standard estimate of pseudo-critical temperature